# Split Gibbs Discrete Diffusion Posterior Sampling

### v1

@@ A.3 Hyperparameters
Appendix B Discussions
B.1 Sample Efficiency
We discuss the sampling efficiency of our method. The time cost of diffusion posterior sampling methods is closely related to the number of neural function evaluations (NFE) of the diffusion model. We present the detailed configurations of SG-DPS with different NFE budgets in Table 6. We report both the total NFE for the overall computation cost and the sequential NFE for the computation cost when each method is fully parallelized.
Configuration ##\## Iterations Prior sampling steps NFE Runtime (sec/sample)
DPS $100$ $102400/100$ 820
SVDD-PM ($M=20$) $128$ $2560/128$ 16
SG-DPS-$50$ $25$ $2$ $50$ 4
SG-DPS-$100$ $25$ $4$ $100$ 6
SG-DPS-$200$ $40$ $5$ $200$ 8
SG-DPS-$400$ $40$ $10$ $400$ 10
SG-DPS-$1$k $50$ $20$ $1000$ 13
SG-DPS-$2$k $100$ $20$ $0020$
Table 6: Configurations of SG-DPS and baseline methods. For DPS and SVDD-PM, we calculate both the total NFE and the sequential NFE. The numbers are collected from the MNIST XOR task as an example. The runtime is amortized on a batch size of 10 samples.

### v2

@@ C.3 Hyperparameters
Synthetic DNA design MNIST XOR MNIST AND Music infilling
Metropolis-Hastings $T$ 10 200 2000 5000 5000
SGDD iterations $K$ 10 50 50 100 100
Euler sampler $H$ 20 20 20 20 20
Sequence length $D$ 2/5/10 200 1024 1024 256
Vocab size $N$ 50 4 2 2 129
Appendix D Additional Experimental Results
D.1 Solving image restoration problems in discrete latent space.
To demonstrate the performance of SGDD on higher-dimensional datasets, we conduct an additional experiment on FFHQ 256 with a vector-quantized diffusion model. Specifically, we use a pretrained VQVAE model from [13] and train a discrete diffusion model in the discrete latent space. We test our algorithm on the super-resolution task ($\times$4) using 100 images from the test dataset. While existing methods fail to solve this high-dimensional task with a complicated forward model that includes a latent space decoder (the cost of DPS is unaffordable), our method is able to reconstruct the underlying images with high PSNR values.
While the latent codes of VQVAE lie in a discrete-state space, they are embedded in an underlying Euclidean space, which makes this problem different from solving inverse problems in categorical distributions without ordinal information. A potential way to speed up and boost the performance of SGDD in solving image restoration problems with VQ-diffusion is to replace the Metropolis-Hastings in discrete latent codes with Langevin dynamics in the continuous embedding space of latent codes. We leave the exploration of this idea as a future direction.
Table 8: Quantitative results for the super-resolution ($\times$4) task on FFHQ 256 dataset.
PSNR $\uparrow$ LPIPS $\downarrow$
SVDD-PM $12.08_{\pm 0.39}$ $0.594_{\pm 0.061}$
SGDD $\textbf{25.85}_{\pm 0.39}$ $\textbf{0.288}_{\pm 0.039}$
Table 9: Class-conditioned generation on MNIST dataset.
[…]
SMC $39.13$ $45.0$
SGDD 14.59 98.5
D.2 More results for inverse problems on discretized image.
Qualitative results. Here we present the visualization of the reconstructed samples of MNIST from their XOR measurements.
Figure 7: Sampling results of the XOR task on the discretized MNIST dataset.  SGDD faithfully recovers the structural information of the ground truth signal.
Approximate distribution divergence on class-conditioned generation. Since a classifier on MNIST data can also be viewed as a forward model, SGDD can also be directly applied to sample digit-conditioned samples. Moreover, while the posterior distribution for a general inverse problem is inaccessible, we can approximate the posterior distribution of class-conditioned generation by a digit-specific prior distribution.
Specifically, we define the posterior distribution as proportional to $p({\mathbf{x}})p(-\beta\ell({\bm{G}}({\mathbf{x}}),{\mathbf{y}}))$, where $\ell$ is the cross-entropy loss between the predicted class probability and target class label ${\mathbf{y}}$. We let ${\mathbf{y}}=0$ to guide diffusion models to generate digits 00, and extract all digits 00 in the dataset as the approximated true posterior distribution. We train another classifier with a different random seed to evaluate the generated samples. To measure the divergence of the generated samples from the posterior distribution, we extract the features of MNIST before the last layer of a CNN classification network, and compute the multi-scale intrinsic distance (MSID) score.
Experimental results in Table 9 show that SGDD effectively samples from the true posterior distribution, as it achieves a low MSID score and a high class accuracy of $98.5\%$.
D.3 More experimental results on the synthetic dataset
Appendix E Discussions
E.1 Limitations and future directions
Our method is built on discrete diffusion models with a uniform transition matrix. It is unclear how our method can be applied to other types of discrete diffusion models, such as masked diffusion models. Moreover, our method calls the forward model (for inverse problems) or the reward function multiple times while running Metropolis-Hastings in likelihood steps, which potentially increases the computational overhead when the forward model or reward function is complicated. We leave these issues as possible future extensions.
E.2 Boarder Impacts
We anticipate that SGDD offers a clean and principled framework for posterior sampling using discrete diffusion models. SGDD tackles these problems by leveraging a discrete diffusion model as a general denoiser in the finite space, which models the prior distribution with rich data. However, we note that it could be the case that SGDD will generate biased samples, if the discrete diffusion model is itself trained on biased data. Therefore, caution should be exercised when using SGDD in certain sensitive scenarios.